OpenStudy (anonymous):

Find the domain of the function. Express your answer in interval notation. -=----->>

OpenStudy (anonymous):

\[f(x)=\sqrt{3x+1}\]

OpenStudy (anonymous):

\[(-1,+\infty)\]

OpenStudy (anonymous):

3x+1>=0 3x>=-1, x>=-1/3 so [-1/3, +infinity)

OpenStudy (anonymous):

x has to be greater than or equal to 1/3

OpenStudy (anonymous):

function has to be 0 or positive